OpenStudy (anonymous):

please help equations and inequalities .2x+.4y=-4 .4x+.2y=-2

OpenStudy (smartanne):

I can help

OpenStudy (smartanne):

Do you know how to solve system of equations?

OpenStudy (anonymous):

thank you smartanne! how do you solve?

OpenStudy (anonymous):

kind of. having trouble with them

OpenStudy (smartanne):

So first take one of the equations and solve for x

OpenStudy (anonymous):

the LCD for the iliminating is confusing me

OpenStudy (smartanne):

.2x+.4y=-4 .2x=-4-.4y x=-4-.4y/.2

OpenStudy (smartanne):

Simplify that equation in that plug it in (for x) to the second equation and you should get the value of y

OpenStudy (smartanne):

Plug in what you get for y in the first equation that you used and you will get the value of x.

OpenStudy (smartanne):

Medal if this helped :) Message me if you have any more questions..also fan me I'm on twice a day

OpenStudy (gabbyalicorn):

0.2x+0.4y=−4;0.4x+0.2y=−2 Step 1: Solve 0.2x+0.4y=−4for x: 0.2x+0.4y+−0.4y=−4+−0.4y(Add -0.4y to both sides) 0.2x=−0.4y−4 0.2x/0.2=−0.4y−4/0.2(Divide both sides by 0.2) x=−2y−20 Step 2: Substitute−2y−20forxin0.4x+0.2y=−2: 0.4x+0.2y=−2 0.4(−2y−20)+0.2y=−2 −0.6y−8=−2(Simplify both sides of the equation) −0.6y−8+8=−2+8(Add 8 to both sides) −0.6y=6 −0.6y/−0.6=6/−0.6(Divide both sides by -0.6) y=−10 Step 3: Substitute−10 for y in x=−2y−20: x=−2y−20 x=(−2)(−10)−20 x=0(Simplify both sides of the equation) And here's your answer: x=0 and y=−10 That's what I got. :/